# Caiqi Zhang1, Fangyu Liu1, Marco Basaldella2, Nigel Collier1

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capability in a variety of NLP tasks. However, LLMs are also prone to generate nonfactual content. Uncertainty Quantification (UQ) is pivotal in enhancing our understanding of a model’s confidence on its generation, thereby aiding in the mitigation of nonfactual outputs. Existing research on UQ predominantly targets short text generation, typically yielding brief, word-limited responses. However, real-world applications frequently necessitate much longer responses. Our study first highlights the limitations of current UQ methods in handling long text generation. We then introduce Luq and its two variations, a series of novel sampling-based UQ approaches specifically designed for long text. Our findings reveal that Luq outperforms existing baseline methods in correlating with the model’s factuality scores (negative coefficient of -0.85 observed for Gemini Pro). To further improve the factuality of LLM responses, we propose Luq-Ensemble, a method that ensembles responses from multiple models and selects the response with the lowest uncertainty. The ensembling method greatly improves the response factuality upon the best standalone LLM.
@@ 1 Introduction
Although Uncertainty Quantification (UQ) is a well-researched area in machine learning (Gawlikowski et al., 2023), its application in the context of LLMs remains under-explored. One primary limitation is that previous studies on UQ mostly require access to a model’s internal states (e.g. logits) (Murray and Chiang, 2018; Kuhn et al., 2022; Vazhentsev et al., 2023; Duan et al., 2023). However, many best-performing LLMs, such as GPT-4 (OpenAI, 2023), Gemini 1.0 Pro (Gemini Team, 2023), and Claude 2.1 (Anthropic, 2023), are closed-source and only accessible via API calls. This limits the ability to directly analyze their internal processes. Another challenge is that existing research on modeling uncertainty predominantly focuses on short responses, typically less than 10 words in length (Kuhn et al., 2022; Duan et al., 2023; Lin et al., 2023). This is in stark contrast to the more common use cases of LLMs, where responses to queries often far exceed this length, sometimes reaching hundreds of words. Such disparity points to a need for new UQ methods tailored for long-form text generated by LLMs. Therefore, in this study we aim to answer the following research questions: RQ1: Are existing UQ methods still effective in the context of long-text generation? RQ2: If not, how can we effectively quantify LLMs’ uncertainty for long-form answers? RQ3: In what ways can uncertainty scores be utilized to enhance the factuality of model outputs?
@@
  • We first highlight the limitations of existing UQ methods for long text generation and then propose Luq (Long-text Uncertainty Quantification; pronounced as “luck”), a novel UQ method that computes sentence-level consistency in long text scenarios.
@@ 2.1 Uncertainty and Confidence
Currently, a formal and universally accepted definition of uncertainty levels in language generation tasks remains elusive. Common practice in existing literature measures uncertainty through the entropy of predictions, akin to approaches in classification tasks (Kuhn et al., 2022; Lin et al., 2023). Predictive entropy is formally expressed as:
@@
In classification tasks, confidence for a specific prediction $y$ is quantified using the predicted probability, represented as $\hat{p}(Y=y\mid x)$ (Geifman and El-Yaniv, 2017; Hendrycks and Gimpel, 2016). Similarly, in the context of NLG, the confidence score for a given response $\mathbf{r}$ is represented by the joint probability of the tokens in the response:
@@ 2.2 Uncertainty for Long Text Generation
In our study, we adopt a more flexible approach to defining uncertainty and confidence in long text generation. Similar to Huang et al. (2024), we focus on the ability of UQ methods to effectively rank responses, differentiating between correct and incorrect predictions. This approach also aligns with the concept of relative confidence as discussed by Geng et al. (2023). Our objective diverges from the orthogonal research direction about models’ calibration, which requires models to precisely reflect their true accuracy in practical scenarios (Lin et al., 2023). We argue that while short-answer questions may be straightforwardly assessed using metrics such as accuracy or exact match, these standards are often unrealistic for long text generation, given the complexities of real-life probabilities.
@@ 3 Luq
Following the generation of $n$ responses, traditional UQ methods for short text commonly calculate the pairwise similarity among the responses (Kuhn et al., 2022; Lin et al., 2023). These pairwise similarity scores indicate the consistency between a pair of responses and play a vital role in subsequent uncertainty estimation. However, answers to certain questions such as “Give me an introduction of …” and “Tell me something about …” may extend to hundreds of words. Longer text leads to an unexpected high similarity across all response pairs when applying previous methods. To address this issue and achieve a more nuanced similarity assessment, we propose the Luq uncertainty measurement with sentence-level similarity computation. Inspired by the hallucination detection method in Manakul et al. (2023), we split each response to sentences, and check whether each sentence can be supported by other samples.
@@
Unlike Kuhn et al. (2022)’s method of applying an off-the-shelf DeBERTa model, we apply the DeBERTa-v3-large model (He et al., 2022), fine-tuned on the MultiNLI (Williams et al., 2018) dataset. This choice is due to our input being a concatenation of short hypothesis (sentence $s$) and a comparatively longer premises (reference response $r^{\prime}$). The format of our input aligns with the task in MultiNLI dataset, ensuring an effective assessment of consistency among the responses.
@@ 4.1 Dataset, Metric, and LLM Selection
Baselines for UQ. We use the following black-box UQ methods as baselines: Lexical similarity (LexSim) (Fomicheva et al., 2020), Number of semantic sets (NumSets) (Lin et al., 2023), Sum of eigenvalues of the graph Laplacian (EigV) (Lin et al., 2023), Degree matrix (Deg) (Lin et al., 2023), Eccentricity (Ecc) (Lin et al., 2023), SelfCheckNLI (SCN) Manakul et al. (2023). We also include three white-box methods for comparison: Maximum Sequence Probability (MSP), Monte Carlo Sequence Entropy (MCSE) (Malinin and Gales, 2020), and Semantic Entropy (SE) (Kuhn et al., 2022). More details can be found in Appendix E.
@@ 4.2 Uncertainty Quantification Results
Luq for GPT-4. We also observe that Luq is better suited for models with relatively lower factuality and a lack of self-expressiveness regarding uncertainty. For models with high factuality capabilities, such as GPT-4, Luq only demonstrates a moderate correlation with factuality scores. As shown in Table 2, among all models, GPT-4 exhibits the highest overall factuality scores and the lowest average uncertainty scores. Figure 2(a) also shows that the data points of GPT-4 are tightly clustered with only few instances of high uncertainty. This is because GPT-4 tends to abstain from answering questions more often compared to other models, highlighting improved uncertainty self-detection w.r.t. ChatGPT-3.5. This observation does not influence the effectiveness of our method, as in real life models with lower factuality and unable to express uncertainty are in greater need of external uncertainty measurements.
@@
Table 4: Results of different ensemble strategies on FActScore-Bio (expressed as percentage). The Answer Distribution (AD) metric indicates the percentage of final answers generated by each component model.
@@ 4.3 Luq-Ensemble
Given the variance in training corpus, different LLMs may possess varying levels of knowledge for a specific question. After obtaining outputs from different LLMs, the challenge now is to choose the best one without fact-checking each answer (which is both time-consuming and costly). Utilizing the Luq uncertainty score as a reliable indicator of factuality, we enhance overall performance through an ensemble approach. In this method, the model exhibiting the lowest Luq score for a given question is chosen as the final answer. Experimental results (Table 4) affirm the superiority of the Luq-Ensemble over its constituent counterparts.
@@
Ensembling does not guarantee better performance. While ensembling always reduces uncertainty scores (as we select the model with the least uncertainty), it does not necessarily improve factuality scores. Ensembling LLMs according to poor UQ methods may result in overall performance that is worse than that of its individual components. Table 10 in Appendix G compares the effectiveness of using LUQ as the ensemble indicator with other methods. The results indicate that ensembling does not inherently enhance performance. For example, with a certain UQ method like Ecc, the ensemble factuality score can be lower than that of its best-performing component (47.2 vs 43.3). In contrast, using Luq as the ensembling indicator yields the best overall performance.
@@ 4.4 Selective Question Answering
From Table 2, it is observed that while GPT-4 opts not to respond to some queries, other models generally attempt to answer all questions. The limited refusal by Gemini 1.0 Pro primarily stems from considerations of sensitive content and regulatory constraints, rather than uncertainty. Therefore, we investigate the application of the Luq score to equip these models with the capability for selective question answering—that is, to enable them to decline responses when uncertain. Contrary to the traditional aim of responding correctly to every question, the objective in a selective question answering framework is to preserve accuracy while maximizing the number of questions answered (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023).
@@ 5 Related Work
UQ in LLMs. In contrast to discriminative models, which readily provide probability scores for specific categories, uncertainty estimation in generative LLMs presents unique challenges: (1) There is an exponential increase in the output space as sentence length grows, rendering the evaluation of all possible predictions impractical (Geng et al., 2023; Wang et al., 2023). (2) The significance of semantic nuances and their inherent uncertainties, which diverges from the fixed category labels typical of discriminative models, complicates matters further (Kuhn et al., 2022). Generally, UQ methods for LLMs can be categorized based on the accessibility of the model’s internal states, distinguishing between black-box and white-box approaches. White-box LLMs often rely on logit-based evaluations, assessing sentence uncertainty through token-level probabilities or entropy (Murray and Chiang, 2018; Kuhn et al., 2022; Vazhentsev et al., 2023; Duan et al., 2023).
However, as access to LLMs increasingly relies on API calls, research has pivoted towards black-box methods. These can be further categorized into: (i) verbalized methods, which prompt LLMs to articulate their uncertainty in the output, using phrases like “I am sure" or “I do not know" (Mielke et al., 2022). Nonetheless, a practical mismatch between the expressed and actual uncertainty levels has been noted (Lin et al., 2022; Xiong et al., 2023). Xiong et al. (2023) highlight that LLMs often display excessive confidence when verbalizing their certainty. (ii) Consistency-based (sampling-based) estimation premises on the assumption that increased uncertainty in a model corresponds to greater diversity in its outputs, frequently resulting in hallucinatory outputs (Manakul et al., 2023; Lin et al., 2023). Our proposed method, Luq, follows this consistency-based approach. There are also efforts on integrating verbalized methods with consistency-based approaches (Xiong et al., 2023; Rivera et al., 2024).
@@ 6 Conclusion
The limitations of this study include the following: (1) A primary challenge in studying uncertainty quantification for long text generation lies in the difficulty of evaluating the generated text. Unlike classification tasks and short-answer QA, there is no straightforward metric for assessing the quality of generated text. In this study, we employ the factuality score as the primary evaluation metric, thereby leaving other text aspects, such as coherence, cohesion, and creativity, under-explored. Future work could investigate uncertainty scores using more comprehensive evaluation metrics. (2) Fact-checking for long text generation is costly, especially when involving human annotation. While we utilized FActScore as an automated evaluation tool, it still requires numerous API calls to fact-check each piece of information. The development of more cost-effective evaluation metrics or datasets could allow for the expansion of the experimental scope. (3) We discovered that temperature plays a crucial role in measuring uncertainty. A common limitation among all consistency-based uncertainty quantification methods is their effectiveness only at high temperatures. In practice, if some scenarios necessitate a fixed low temperature, these consistency-based UQ methods may not perform effectively. (4) In this study, we do not investigate the performance of UQ methods under ambiguous and unanswerable questions, such as ASQA (Stelmakh et al., 2022) and SelfAware (Yin et al., 2023). Previous uncertainty metrics for short-answer questions are tested on answerable questions with clear intentions. This is because clearly defined questions with definite answers provide a straightforward framework for evaluating model accuracy. In contrast, unanswerable or ambiguous questions lack clear ground truths, complicating the assessment of uncertainty estimates. We advocate researchers to explore more in this area in the future.
@@ Appendix D Experiment Setup
For GPT-4 and GPT-3.5, we use the OpenAI API, with specific version gpt-4-turbo-0125-preview and gpt-3.5-turbo-0613. For Gemini 1.0 Pro, we call the API for developers. For Yi-34B-Chat, Tulu-2-70B (tulu-2-dpo-70b), and Vicuna-33B (vicuna-33b-v1.3), we use them off-the-shelf and only for inference. We run our uncertainty measurement experiments on A100-SXM-80GB GPUs. For our experiments, we use the following prompt:
@@ Appendix I Ablation Study
Previous research on short answer generation (Kuhn et al., 2022; Lin et al., 2023) has demonstrated that an increase in the number of samples correlates with enhanced performance. We investigate whether it also applies to long-text generation and find that with more samples, Luq shows better performance and PCC scores, which corroborates with previous observations in short-text generation, as depicted in Figure5. Providing a greater number of samples enables the NLI process to predict sentence factuality with higher accuracy. However, a notable drawback of increasing the sample size is the associated rise in computational costs.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capability in a variety of NLP tasks. However, LLMs are also prone to generate nonfactual content. Uncertainty Quantification (UQ) is pivotal in enhancing our understanding of a model’s confidence on its generation, thereby aiding in the mitigation of nonfactual outputs. Existing research on UQ predominantly targets short text generation, typically yielding brief, word-limited responses. However, real-world applications frequently necessitate much longer responses. Our study first highlights the limitations of current UQ methods in handling long text generation. We then introduce Luq with its two variations: Luq-Atomic and Luq-Pair, a series of novel sampling-based UQ approaches specifically designed for long text. Our findings reveal that Luq outperforms existing baseline methods in correlating with the model’s factuality scores (negative coefficient of -0.85 observed for Gemini Pro). To further improve the factuality of LLM responses, we propose Luq-Ensemble, a method that ensembles responses from multiple models and selects the response with the lowest uncertainty. The ensembling method greatly improves the response factuality upon the best standalone LLM. [^1] https://github.com/caiqizh/LUQ
@@ 1 Introduction
Although Uncertainty Quantification (UQ) is a well-researched area in machine learning (Gawlikowski et al., 2023), its application in the context of LLMs remains under-explored. One primary limitation is that previous studies on UQ mostly require access to a model’s internal states (e.g., logits) (Murray and Chiang, 2018; Kuhn et al., 2023; Vazhentsev et al., 2023; Duan et al., 2023). However, many best-performing LLMs, such as GPT-4 (OpenAI, 2023), Gemini 1.0 Pro (Gemini Team, 2023), and Claude 2.1 (Anthropic, 2023), are closed-source and only accessible via API calls. This limits the ability to directly analyze their internal processes. Another challenge is that existing research on modeling uncertainty predominantly focuses on short responses, typically less than 10 words in length (Kuhn et al., 2023; Duan et al., 2023; Lin et al., 2023). This is in stark contrast to the more common use cases of LLMs, where responses to queries often far exceed this length, sometimes reaching hundreds of words. Such disparity points to a need for new UQ methods tailored for long-form text generated by LLMs. Therefore, in this study we aim to answer the following research questions: RQ1: Are existing UQ methods still effective in the context of long-text generation? RQ2: If not, how can we effectively quantify LLMs’ uncertainty for long-form answers? RQ3: In what ways can uncertainty scores be utilized to enhance the factuality of model outputs?
@@
  • We first highlight the limitations of existing UQ methods for long text generation and then propose Luq (Long-text Uncertainty Quantification; pronounced as [Uncaptioned image]luck), a novel UQ method that computes sentence-level consistency in long text scenarios.
@@ 2.1 Uncertainty and Confidence
Currently, a formal and universally accepted definition of uncertainty levels in language generation tasks remains elusive. Common practice in existing literature measures uncertainty through the entropy of predictions, akin to approaches in classification tasks (Kuhn et al., 2023; Lin et al., 2023). Predictive entropy is formally expressed as:
@@
In classification tasks, confidence for a specific prediction $y$ is quantified using the predicted probability, represented as $\hat{p}(Y=y\mid x)$ (Geifman and El-Yaniv, 2017; Hendrycks and Gimpel, 2017). Similarly, in the context of NLG, the confidence score for a given response $\mathbf{r}$ is represented by the joint probability of the tokens in the response:
@@ 2.2 Uncertainty for Long Text Generation
In our study, we adopt a more flexible approach to defining uncertainty and confidence in long text generation. Similar to Huang et al. (2024a), we focus on the ability of UQ methods to effectively rank responses, differentiating between correct and incorrect predictions. This approach also aligns with the concept of relative confidence as discussed by Geng et al. (2023). Our objective diverges from the orthogonal research direction about models’ calibration, which requires models to precisely reflect their true accuracy in practical scenarios (Lin et al., 2023). We argue that while short-answer questions may be straightforwardly assessed using metrics such as accuracy or exact match, these standards are often unrealistic for long text generation, given the complexities of real-life probabilities.
@@ 3 Luq
Following the generation of $n$ responses, traditional UQ methods for short text commonly calculate the pairwise similarity among the responses (Kuhn et al., 2023; Lin et al., 2023). These pairwise similarity scores indicate the consistency between a pair of responses and play a vital role in subsequent uncertainty estimation. However, answers to certain questions such as “Give me an introduction of …” and “Tell me something about …” may extend to hundreds of words. Longer text leads to an unexpected high similarity across all response pairs when applying previous methods. To address this issue and achieve a more nuanced similarity assessment, we propose the Luq uncertainty measurement with sentence-level similarity computation. Inspired by the hallucination detection method in Manakul et al. (2023), we split each response to sentences, and check whether each sentence can be supported by other samples.
@@
Unlike Kuhn et al. (2023)’s method of applying an off-the-shelf DeBERTa model, we apply the DeBERTa-v3-large model (He et al., 2023), fine-tuned on the MultiNLI (Williams et al., 2018) dataset. This choice is due to our input being a concatenation of short hypothesis (sentence $s$) and a comparatively longer premises (reference response $r^{\prime}$). The format of our input aligns with the task in MultiNLI dataset, ensuring an effective assessment of consistency among the responses.
@@ 4.1 Dataset, Metric, and LLM Selection
Baselines for UQ. We use the following black-box UQ methods as baselines: Lexical similarity (LexSim) (Fomicheva et al., 2020), Number of semantic sets (NumSets) (Lin et al., 2023), Sum of eigenvalues of the graph Laplacian (EigV) (Lin et al., 2023), Degree matrix (Deg) (Lin et al., 2023), Eccentricity (Ecc) (Lin et al., 2023), SelfCheckNLI (SCN) Manakul et al. (2023). We also include three white-box methods for comparison: Maximum Sequence Probability (MSP), Monte Carlo Sequence Entropy (MCSE) (Malinin and Gales, 2021), and Semantic Entropy (SE) (Kuhn et al., 2023). More details can be found in Appendix E.
@@ 4.2 Uncertainty Quantification Results
Luq for GPT-4. We also observe that Luq is better suited for models with relatively lower factuality and a lack of self-expressiveness regarding uncertainty. For models with high factuality capabilities, such as GPT-4, Luq only demonstrates a moderate correlation with factuality scores. As shown in Table 2, among all models, GPT-4 exhibits the highest overall factuality scores and the lowest average uncertainty scores. Figure 2(a) also shows that the data points of GPT-4 are tightly clustered with only few instances of high uncertainty. This is because GPT-4 tends to abstain from answering questions more often compared to other models, highlighting improved uncertainty self-detection. However, this observation does not influence the effectiveness of our method, as in real life models with lower factuality and unable to express uncertainty are in greater need of external uncertainty measurements.
@@
Table 4: Results of different ensemble strategies on FActScore-Bio (expressed as percentage). The Answer Distribution (AD) indicates the percentage of final answers generated by each component model.
@@ 4.3 Luq-Ensemble
Given the variance in training corpus, different LLMs may possess varying levels of knowledge for a specific question. After obtaining outputs from different LLMs, the challenge now is to choose the best one without fact-checking each answer (which is both time-consuming and costly (Guo et al., 2022; Zhang et al., 2024)). Utilizing the Luq uncertainty score as a reliable indicator of factuality, we enhance overall performance through an ensemble approach. In this method, the model exhibiting the lowest Luq score for a given question is chosen as the final answer. Experimental results (Table 4) affirm the superiority of the Luq-Ensemble over its constituent counterparts.
@@
Ensembling does not guarantee better performance. While ensembling always reduces uncertainty scores (as we select the model with the least uncertainty), it does not necessarily improve factuality scores. Ensembling LLMs according to poor UQ methods may result in overall performance that is worse than that of its individual components. Table 10 in Appendix G compares the effectiveness of using LUQ as the ensemble indicator with other methods. The results indicate that ensembling does not inherently enhance performance. For example, with UQ method Ecc, the ensemble factuality score can be lower than that of its best-performing component (47.2% vs 43.3%). In contrast, using Luq as the ensembling indicator yields the best overall performance.
@@ 4.4 Selective Question Answering
From Table 2, it is observed that while GPT-4 opts not to respond to some queries, other models generally attempt to answer all questions. The limited refusal by Gemini 1.0 Pro primarily stems from considerations of sensitive content and regulatory constraints, rather than uncertainty. Therefore, we investigate the application of the Luq score to equip these models with the capability for selective question answering—that is, to enable them to decline responses when uncertain. Contrary to the traditional aim of responding correctly to every question, the objective in a selective question answering framework is to preserve accuracy while maximizing the number of questions answered (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023; Dong et al., 2024).
@@ 5 Related Work
UQ in LLMs. In contrast to discriminative models, which readily provide probability scores for specific categories, uncertainty estimation in generative LLMs presents unique challenges: (1) There is an exponential increase in the output space as sentence length grows, rendering the evaluation of all possible predictions impractical (Geng et al., 2023; Wang et al., 2023). (2) The significance of semantic nuances and their inherent uncertainties, which diverges from the fixed category labels typical of discriminative models, complicates matters further (Kuhn et al., 2023). Generally, UQ methods for LLMs can be categorized based on the accessibility of the model’s internal states, distinguishing between black-box and white-box approaches. White-box LLMs often rely on logit-based evaluations, assessing sentence uncertainty through token-level probabilities or entropy (Murray and Chiang, 2018; Kuhn et al., 2023; Vazhentsev et al., 2023; Duan et al., 2023).
However, as access to LLMs increasingly relies on API calls, research has pivoted towards black-box methods. These can be further categorized into: (i) verbalized methods, which prompt LLMs to articulate their uncertainty in the output, using phrases like “I am sure" or “I do not know" (Mielke et al., 2022). Nonetheless, a practical mismatch between the expressed and actual uncertainty levels has been noted (Lin et al., 2022; Xiong et al., 2023). Xiong et al. (2023) highlight that LLMs often display excessive confidence when verbalizing their certainty. (ii) Consistency-based (sampling-based) estimation premises on the assumption that increased uncertainty in a model corresponds to greater diversity in its outputs, frequently resulting in hallucinatory outputs (Manakul et al., 2023; Lin et al., 2023). Our proposed method, Luq, follows this consistency-based approach. There are also efforts on integrating verbalized methods with consistency-based approaches (Xiong et al., 2023; Rivera et al., 2024). Understanding uncertainty in LLMs can enhance in-context learning (Zhou et al., 2023; Li et al., 2023), selective question answering (Yang et al., 2023), LLM cascading (Huang et al., 2024b), adaptive retrieval (Ding et al., 2024), language agents (Han et al., 2024), and model self-refinement (Yao et al., 2024; Chen et al., 2024).
@@ 6 Conclusion
The limitations of this study include the following: (1) A primary challenge in studying uncertainty quantification for long text generation lies in the difficulty of evaluating the generated text. Unlike classification tasks and short-answer QA, there is no straightforward metric for assessing the quality of generated text. In this study, we employ the factuality score as the primary evaluation metric, thereby leaving other text aspects, such as coherence, cohesion, and creativity, under-explored. Future work could investigate uncertainty scores using more comprehensive evaluation metrics. (2) In this study, we do not investigate the performance of UQ methods under ambiguous and unanswerable questions, such as ASQA (Stelmakh et al., 2022) and SelfAware (Yin et al., 2023). Previous uncertainty metrics for short-answer questions are tested on answerable questions with clear intentions. This is because clearly defined questions with definite answers provide a straightforward framework for evaluating model accuracy. In contrast, unanswerable or ambiguous questions lack clear ground truths, complicating the assessment of uncertainty estimates. We advocate researchers to explore more in this area in the future. (3) We focus on assessing the overall uncertainty of a model, rather than model uncertainty on individual instances. The relative uncertainty equation in Section 2.2 represents an ideal scenario. If a model learns a significant amount of non-factual data over factual data for a particular entity/instance, the aforementioned equation can be inaccurate for that case. Future work could investigate the causes of this special case and develop strategies to address it during the pre-training stage.
@@
Acknowledgments
Caiqi Zhang is supported by an Amazon Studentship. We thank Chengzu Li, Yulong Chen, and the reviewers for their valuable feedback on this paper.
@@ Appendix D Experiment Setup
For GPT-4 and GPT-3.5, we use the OpenAI API, with specific version gpt-4-turbo-0125-preview and gpt-3.5-turbo-0613. For Gemini 1.0 Pro, we call the API for developers. For Yi-34B-Chat, Tulu-2-70B (tulu-2-dpo-70b), and Vicuna-33B (vicuna-33b-v1.3), we use them off-the-shelf and only for inference. The temperature is set to 0.7. We run our uncertainty measurement experiments on A100-SXM-80GB GPUs. For our experiments, we use the following prompt:
@@ Appendix I Ablation Study
Previous research on short answer generation (Kuhn et al., 2023; Lin et al., 2023) has demonstrated that an increase in the number of samples correlates with enhanced performance. We investigate whether it also applies to long-text generation and find that with more samples, Luq shows better performance and PCC scores, which corroborates with previous observations in short-text generation, as depicted in Figure5. Providing a greater number of samples enables the NLI process to predict sentence factuality with higher accuracy. However, a notable drawback of increasing the sample size is the associated rise in computational costs.